OpenStudy (stormoshawty):

Solve this equation using elimination. 3x + 3y =27 x - 3y = -11

OpenStudy (cloverracer):

Algebra I Connections Academy?

OpenStudy (stormoshawty):

Yes.

OpenStudy (shamil98):

add them together

OpenStudy (anonymous):

x=4 y=5

OpenStudy (anonymous):

add the two equations together so u get 6x=16 divide both sides by 6 x=16/6 simplify x=8/3

OpenStudy (shamil98):

@ngott3 that's wrong

OpenStudy (shamil98):

3x + 3y =27 x - 3y = -11 adding them you get 4x = 16 storm can you do it from here?

OpenStudy (shamil98):

once you find x sub it back into x - 3y = -11 to get your value for y.

OpenStudy (anonymous):

omg what am i doing -.-" thanx

OpenStudy (anonymous):

i really shouldn't be up at 2:03 am in the morning

OpenStudy (stormoshawty):

Y=5?

OpenStudy (shamil98):

4 - 3y = -11 -3y = -15 y = 5 yep good work

OpenStudy (anonymous):

yes and x=4